# Crossed products by compact group actions with the weak tracial Rokhlin property

### v1

@@ 7. A totally disconnected compact group action
By [33, Theorem 1.10.16], $A\cong\mathcal{Z}$.
@@
This follows from [27, Example 5.10] (see also [1, Theorem 3.10]) and Proposition 6.3. ∎
@@
Since $A\cong\mathcal{Z}$, we let $\delta_{1}=d_{\tau}(x)>0$ where $\tau$ is the unique tracial state on $A$.
By [27, Example 5.10], the action $\gamma$ has the weak tracial Rokhlin property. Therefore, by [12, Lemma 3.3], $B\rtimes_{\gamma}S_{2}$ is simple, so by Theorem 4.5, $B^{\gamma}$ is simple. It is easy to verify that $A^{\alpha}$ can be identified with $\bigotimes_{m=1}^{\infty}B_{m}^{\gamma}$, so it is simple. Also, $B$ has strict comparison, so $B^{\gamma}$ does too, and thus $A^{\alpha}$ does too (see [11, Proposition 5.16.1(4)]). Since $A^{\alpha}$ is unital, stably finite, and exact, $\mathrm{T}(A^{\alpha})$ is nonempty. We let $\delta_{2}=\inf_{\tau\in\mathrm{T}(A^{\alpha})}d_{\tau}(y)>0$.
Since $\cup_{n=1}^{\infty}A_{n}$ is dense in $A$, there exist $N_{1}\in\mathbb{N}$ and a finite set $F_{0}\subseteq A_{N_{1}}\subseteq A$ such that for each $a\in F$ there exists $b\in F_{0}$ satisfying $\|a-b\|<\frac{\varepsilon}{8}$, and $\|b\|\leq 1$ for all $b\in F_{0}$. We can assume there exist $c_{ij}\in B$ for $i=1,2,\dots,N_{1}$ and $j=1,2,\dots,k$ such that $F_{0}=\{c_{1j}\otimes c_{2j}\otimes\dots\otimes c_{N_{1}j}\colon 1\leq j\leq k\}$. Similarly, there exist $N_{2}\in\mathbb{N}$ and a finite set $S_{0}\subseteq C(G_{N_{2}})\subseteq C(G)$ such that for each $f\in S$ there exists $e\in S_{0}$ satisfying $\|f-e\|<\frac{\varepsilon}{4}$ and $\|e\|\leq 1$ for all $e\in S_{0}$. Let $N=\max(N_{1},N_{2})$. Let $M=\mathrm{card}(G_{N})$, and let $\varepsilon_{0}=\min(\frac{\delta_{1}}{N},\frac{\delta_{2}}{N},\frac{1}{2N},\frac{\varepsilon}{4MN})$.
Since the action $\gamma$ has the weak tracial Rokhlin property. By Lemma 7.4, for each $1\leq i\leq N$, letting $F_{i}=\{c_{ij}\colon 1\leq j\leq k\}$, we obtain a contractive positive element $d^{(i)}\in B^{\gamma}$ and mutually orthogonal contractive positive elements $(d_{g}^{(i)})_{g\in H_{i}}\in B$ such that
@@
    $d_{\tau}(1-d^{(i)})<\varepsilon_{0}$ for all $\tau\in\mathrm{T}(B^{\gamma})$.
@@
It is easy to verify that $d_{g}d_{h}=0$ for $g,h\in G_{N}$ with $g\neq h$, $\alpha_{N_{g}}(d_{h})=d_{gh}$ for $g,h\in G_{N}$, and $\|ad_{g}-d_{g}a\|<N\varepsilon_{0}$ for $a\in F_{0}$ and $g\in G_{N}$. Define $d=\sum_{h\in G_{N}}d_{h}\in A_{N}\subseteq A$. There exists an equivariant completely positive contractive order zero map $\psi_{0}\colon C(G_{N})\to\overline{dA_{N}d}\subseteq\overline{dAd}$ given by $\psi_{0}(f)=\sum_{h\in G_{N}}f(h)d_{h}$ for $f\in C(G_{N})$, so that $d=\psi_{0}(1)$.
@@
Let $K=\prod_{m=N+1}^{\infty}H_{m}$, so $G=G_{N}\times K$. Let $\mu$ be the normalized Haar measure on $K$. Then there is a conditional expectation $P\colon C(G)\to C(G_{N})$, given by $P(f)(h)=\int_{K}f(h,g)d_{\mu}(g)$ for $f\in C(G)$ and $h\in G_{N}$. By identifying $C(G_{N})$ as a subalgebra of $C(G)$, it is invariant under the left translation action $g\mapsto\mathrm{Lt}_{g}$ of $G$. Under this identification, the map $P$ is equivariant. It is easy to verify that $d\in A^{\alpha}$. Therefore $\psi=\psi_{0}\circ P\colon C(G)\to\overline{dAd}$ is an equivariant completely positive contractive map with $\psi(1)=d$.
@@
$\|b-a\|<\frac{\varepsilon}{8},\ \|e_{1}-f_{1}\|<\frac{\varepsilon}{4},\ \mbox{and}\ \|e_{2}-f_{2}\|<\frac{\varepsilon}{4}.$
@@
Since $d=d^{(1)}\otimes d^{(2)}\otimes\dots\otimes d^{(N)}$, letting $\tau$ be the unique tracial state on $A$, by Lemma 7.6, we have
@@
Similarly, for any $\tau\in\mathrm{T}(A^{\alpha})$, since the set of tracial states on the minimal tensor product of C*-algebras is the closed convex hull of the set of tensor products of tracial states on the factor C*-algebras, we may assume $\tau$ is an elementary tensor product. Therefore, we have
@@
Since $d_{\tau}(d)\geq 1-d_{\tau}(1-d)>\frac{1}{2}>d_{\tau}(1-d)$, and $A^{\alpha}$ has strict comparison, we obtain $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}1-d\precsim_{A^{\alpha}}d$.
Since $A\cong\mathcal{Z}$ is stably finite, the condition $\|dxd\|>1-\varepsilon$ is redundant. ∎

### v2

@@ 7. A totally disconnected compact group action
By [34, Theorem 1.10.16], $A\cong\mathcal{Z}$.
@@
This follows from [28, Example 5.10] (see also [1, Theorem 3.10]) and Proposition 6.3. ∎
@@
Since $A\cong\mathcal{Z}$, we denote $\delta_{1}=d_{\tau}(x)>0$ where $\tau$ is the unique tracial state on $A$.
By [28, Example 5.10], the action $\gamma$ has the weak tracial Rokhlin property. Therefore, by [13, Lemma 3.3], $B\rtimes_{\gamma}S_{2}$ is simple, and hence by Theorem 4.5, $B^{\gamma}$ is simple. It is easy to verify that $A^{\alpha}$ can be identified with $\bigotimes_{m=1}^{\infty}B_{m}^{\gamma}$ (see [38, Lemma 4.5]), hence it is simple. Meanwhile, since $B$ has strict comparison, $B^{\gamma}$ also has it, and thus $A^{\alpha}$ does as well (see [12, Proposition 5.16.1(4)]). Since $A^{\alpha}$ is unital, stably finite, and amenable, $\mathrm{T}(A^{\alpha})$ is a nonempty compact set. We denote $\delta_{2}=\inf_{\tau\in\mathrm{T}(A^{\alpha})}d_{\tau}(y)>0$.
Since $\cup_{n=1}^{\infty}A_{n}$ is dense in $A$, there exists $N_{1}\in\mathbb{N}$ and a finite set $F_{0}\subseteq A_{N_{1}}\subseteq A$ such that for each $a\in F$, there exists $b\in F_{0}$ satisfying $\|a-b\|<\frac{\varepsilon}{8}$, and for all $b\in F_{0}$, $\|b\|\leq 1$. We may assume there exist $c_{ij}\in B$ for $i=1,2,\dots,N_{1}$ and $j=1,2,\dots,k$ such that $F_{0}=\{c_{1j}\otimes c_{2j}\otimes\dots\otimes c_{N_{1}j}\colon 1\leq j\leq k\}$. Similarly, there exists $N_{2}\in\mathbb{N}$ and a finite set $S_{0}\subseteq C(G_{N_{2}})\subseteq C(G)$ such that for each $f\in F$, there exists $e\in S_{0}$ satisfying $\|f-e\|<\frac{\varepsilon}{4}$, and for all $e\in S_{0}$, $\|e\|\leq 1$. Let $N=\max(N_{1},N_{2})$. Let $M=\mathrm{card}(G_{N})$, and let $\varepsilon_{0}=\min(\frac{\delta_{1}}{N},\frac{\delta_{2}}{N},\frac{1}{2N},\frac{\varepsilon}{4MN})$.
Since the action $\gamma$ has the weak tracial Rokhlin property, by Lemma 7.4, for each $1\leq i\leq N$, set $F_{i}=\{c_{ij}\colon 1\leq j\leq k\}$, we obtain a contractive positive element $d^{(i)}\in B^{\gamma}$ and mutually orthogonal contractive positive elements $(d_{g}^{(i)})_{g\in H_{i}}\in B$ such that
@@
    For all $\tau\in\mathrm{T}(B^{\gamma})$, $d_{\tau}(1-d^{(i)})<\varepsilon_{0}$.
@@
It is easy to verify that for $g,h\in G_{N}$ with $g\neq h$, $d_{g}d_{h}=0$; for $g,h\in G_{N}$, $\alpha_{N_{g}}(d_{h})=d_{gh}$; and for $a\in F_{0}$ and $g\in G_{N}$, $\|ad_{g}-d_{g}a\|<N\varepsilon_{0}$. Define $d=\sum_{h\in G_{N}}d_{h}\in A_{N}\subseteq A$. There exists an equivariant completely positive contractive order zero map $\psi_{0}\colon C(G_{N})\to\overline{dA_{N}d}\subseteq\overline{dAd}$ given by $\psi_{0}(f)=\sum_{h\in G_{N}}f(h)d_{h}$ for $f\in C(G_{N})$, such that $d=\psi_{0}(1)$.
@@
Let $K=\prod_{m=N+1}^{\infty}H_{m}$, then $G=G_{N}\times K$. Let $\mu$ be the normalized Haar measure on $K$. Then there exists a conditional expectation $P\colon C(G)\to C(G_{N})$, given by $P(f)(h)=\int_{K}f(h,g)d_{\mu}(g)$ for $f\in C(G)$ and $h\in G_{N}$. By identifying $C(G_{N})$ as a subalgebra of $C(G)$, which is invariant under the left translation action $g\mapsto\mathrm{Lt}_{g}$ of the group $G$. Under this identification, the map $P$ is equivariant. It is easy to verify that $d\in A^{\alpha}$. Therefore, $\psi=\psi_{0}\circ P\colon C(G)\to\overline{dAd}$ is an equivariant completely positive contractive map with $\psi(1)=d$, i.e., condition (2) holds.
@@
$\|b-a\|<\frac{\varepsilon}{8},\ \|e_{1}-f_{1}\|<\frac{\varepsilon}{4},\ \text{and}\ \|e_{2}-f_{2}\|<\frac{\varepsilon}{4}.$
@@
Hence condition (1) holds.
For condition (3), since $d=d^{(1)}\otimes d^{(2)}\otimes\dots\otimes d^{(N)}$, let $\tau$ be the unique tracial state on $A$, by Lemma 7.6, we have
@@
For condition (4), similarly, let $\tau\in\mathrm{T}(A^{\alpha})$, for all $\tau\in\mathrm{T}(A^{\alpha})$, since the set of tracial states on the minimal tensor product of C*-algebras is the closed convex hull of the tensor products of tracial states on the component C*-algebras, we may assume $\tau$ is an elementary tensor product. Therefore, we have
@@
Since $d_{\tau}(d)\geq 1-d_{\tau}(1-d)>\frac{1}{2}>d_{\tau}(1-d)$, and $A^{\alpha}$ has strict comparison, we obtain $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}1-d\precsim_{A^{\alpha}}d$. Hence condition (5) holds.
Since $A\cong\mathcal{Z}$ is stably finite, condition (6) ($\|dxd\|>1-\varepsilon$) is redundant. ∎